# Compact and Selective Disclosure for Verifiable Credentials

### v1

@@ V-C Compromised Communication Channel
To prevent $\mathcal{A}$ from passively accessing credential content, secure communication among parties must be established. This can be effectively achieved by employing asymmetric encryption, utilizing the public keys stored in the DID Document of participants. Additionally, tampering with the accumulator value $a$ or $WVC$ would be ineffective, as any modification would require a respective witness, impossible to generate without $sk_{i}$. Any altered or newly generated accumulator value by the adversary $\mathcal{A}$ would not authenticate against the public key $pk_{i}$ since the adversary must know $sk_{i}$ to compute a valid $a$.
@@ VI Performance Evaluation
Experiments were conducted by comparing CSD-JWT to SD-JWT and run on a 14th Gen Intel(R) Core(TM) i9-14900k that features 8 p-cores and 16 e-cores, and 48GB of RAM. For holder-side operations, we utilized the Nordic nRF52840 [46], an RFC7228 class-2 constrained device [23]. The nRF52840 is equipped with a 32-bit ARM Cortex-M4 processor, 1 MB of Flash memory, and 256 KB of RAM. We compared CSD-JWT to SD-JWT as it represents the current state-of-the-art solution and remains the only fully implemented and widely recognized solution available in the field at the time of this work. The experimental findings demonstrate that our mechanism is a valuable solution to enable selective disclosure while maximizing holder privacy. For fairness, all comparisons between CSD-JWT and SD-JWT were conducted using the same claims within the credentials. All experiments were repeated 1,000 times, and the results were averaged.
Figure 2: Issuer overhead for generating VCs.
@@ VI-A Implementation Setup
We implemented a prototype of CSD-JWT [^1] Code will be made available upon acceptance and evaluated its performance while delivering the core functionalities. To effectively build the SD-JWT mechanism, we relied mainly on two libraries: josekit [^2] https://crates.io/crates/josekit, an openssl-based library that provides ease of access to cryptographic signatures, and serde_json [^3] https://crates.io/crates/serde_json, a library that provides API to serialize and deserialize Rust objects to a JSON format. For generating accumulator values and the witnesses associated with claims, we relied on vb-accumulator [^4] https://crates.io/crates/vb-accumulator, developed by DockNetwork, which meets the requirements of [22]. For the underlying ECC, we selected the BN254 curve, which offers the best trade-off between security and parameter length, as accumulator values and witnesses require 256 bits. To evaluate the feasibility of CSD-JWT and SD-JWT on constrained holders, we developed a framework [^5] Code will be made available upon acceptance in C using the nRF Connect SDK and Parson  [^6] https://github.com/kgabis/parson, a lightweight library for managing JSON data structures. Our credentials can include a variable number of claims, each formatted as claim_key:claim_value. For our evaluation, we consider VCs with up to 100 claims, reflecting variations across application domains [47]. For instance, a VC representing an identification document typically contains around 15 claims. In contrast, in healthcare scenarios, the number of claims can vary significantly, ranging from approximately 10 claims, as in the case of the healthcare passport established by many countries for COVID-19 [48], to over 100 in detailed medical reports [49]. Finally, for transparency and reproducibility, we open-sourced the proposed implementations.
Figure 3: Storage requirements (KB) for VCs.
@@
VI-B Credential Issuance
Figure 2 compares the issuance latencies, expressed on a logarithmic scale, of the considered methods. SD-JWT achieves slightly better performance, mainly due to the CPU instructions used by the two mechanisms. Indeed, despite the great improvements in the instruction set of processors regarding finite field operations, hashing algorithms remarkably outperform them. However, it is worth noting that CSD-JWT only requires a few milliseconds even for credentials including 100 claims. Moreover, this does not represent a serious concern since, in real-world scenarios, issuers typically have powerful computing resources to manage the issuance process, significantly reducing these latencies.
VI-C Credential Storage
As shown in Figure 3, CSD-JWT remarkably outperforms SD-JWT even with a low number of claims in the VC. Using identical claims for both mechanisms, the primary difference lies in the size of the credentials and the storage of witnesses versus salts. The storage requirement for our method is much lower because CSD-JWT only contains the accumulator value, which is 256 bits. In contrast, SD-JWT comprises the hashes of the $N$ claims, i.e., $N\times$ 256 bits, making it dependent on the number of attributes. On the other hand, salts used in SD-JWT only require $N\times$ 128 bits, against the $N\times$ 256 bits for witnesses. Therefore, SD-JWT brings an overhead of ($N-2)\times 256$ bits for storing a VC with the same claims. As highlighted in the figure, CSD-JWT saves a significantly larger portion of memory while increasing the number of claims, achieving up to $46\%$ storage reduction. These results demonstrate that CSD-JWT is a valuable solution for selectively disclosing credentials in constrained environments. Moreover, this saving is even more impactful as each holder is expected to maintain multiple credentials.
Figure 5: VP size reduction of CSD-JWT.
@@ VI-D Generating and Verifying Verifiable Presentations
Size of Verifiable Presentation. Figures 4(b) and 4(c) report the results for VCs containing 10 and 100 claims, respectively. As clearly shown, CSD-JWT consistently generates smaller VPs than SD-JWT regardless of the number of disclosed claims. In Figure 4(b), when the holder discloses a single claim included in a VC with 10 claims, CSD-JWT achieves a $67\%$ reduction in the VP size, which decreases to $30\%$ when all the claims included in the VC are disclosed. The savings in the VP size are even more significant when considering VCs including a larger number of claims. Figure 4(c) demonstrates that with CSD-JWT, a VP disclosing 1 claim out of 100 is $93\%$ smaller in size than the same VP produced with SD-JWT. Even when all claims are being disclosed, CSD-JWT still provides a $27\%$ size reduction. Figure 5 illustrates the reduction in VP size by varying the rounded percentage of disclosed claims. These results highlight that CSD-JWT is an effective method to generate VPs with reduced size and network overhead, a feature that is particularly important in scenarios where holders’ devices have limited memory and network capabilities.
Verification of Verifiable Presentations. Finally, we evaluated the time required for a verifier to assess the authenticity of the presentations. Figure 6 shows the verification latencies for claims of different sizes using both SD-JWT and CSD-JWT. The latency for verifying a VP generated with SD-JWT is negligible (up to a few $\mu$s) since the procedure involves the verification of a digital signature, a handful of hashing operations, and a comparison with a list of hashes. In CSD-JWT, verification time is affected by the underlying cryptographic accumulator. However, it remains on the order of a few milliseconds, meeting the response requirement of the Resource Animation Idle Load (RAIL) model proposed by Google [^7] https://web.dev/articles/rail. As a result, the overhead introduced by CSD-JWT is minimal and comparable to that of SD-JWT. Furthermore, verification is performed by verifiers, which are entities typically equipped with more powerful computational resources, making the impact of such overhead even less relevant.
Figure 6: VP verification latency of CSD-JWT.
@@ VI-E Discussion of Results
Credential Issuance. The time required to issue a CSD-JWT is higher than SD-JWT due to the complexity of the operations performed. In particular, the most demanding operation is the witness generation that, as demonstrated in [50], for $w$ witnesses demands at least $\Omega(w)$. However, the issuance time is a few milliseconds and does not represent a concern as trusted authorities acting as issuers may have more powerful computing resources than those used for our experiments. Furthermore, it is worth noting that as these operations are always done on the issuer side, they do not affect at all the computing capabilities of the holder.
Credential Storage. In CSD-JWT, the storage requirement for the VC remains constant, regardless of the number of claims included, as it only contains the accumulator value and minimal data for management. This contrasts with SD-JWT, where credentials contain a salted hash for each claim. Although salts are shorter than the witnesses, CSD-JWT offers a size reduction ratio of $27-93\%$ for VCs with 100 claims, with even greater savings for credentials containing more fields. In constrained environments, resources are shared across different functions. For instance, it is essential that the memory securely stores VCs and firmware, making even KB-level reductions valuable. Additionally, minimizing resource usage improves energy efficiency, a critical consideration in resource-limited settings. Finally, while our evaluation focuses on the overhead reduction for a single VC, identity owners are projected to use multiple VCs by 2026, making the reductions offered by CSD-JWT significantly more remarkable.
Verifiable Presentation. The generation of VPs must be performed directly by holders. This is a logical consequence of the fact that they have to prove ownership over the disclosed information through a secret key. However, creating a VP does not introduce a particular overhead. Similarly to SD-JWT, the holder is only required to create and sign a JWT containing the claims to be disclosed. This makes both CSD-JWT and SD-JWT practical on constrained devices.
On the other hand, we observed remarkable differences in the sizes of the VPs. In SD-JWT, the credential shared during the presentation comprises the hashes of all the claims, thus the size depends on the number of data included. In contrast, CSD-JWT size is constant and only contains the accumulator value. This minimizes the information shared, significantly reducing the network overhead, and offering higher privacy.
@@ VII Related Work
This section reviews the primary selective disclosure mechanisms for VCs, categorizing them based on the underlying cryptographic techniques used to disclose claims. The categories include atomic credentials, hashed values, signature schemes, and Zero-Knowledge Proofs (ZKPs) [51].
@@ VII-A Atomic Credentials
Atomic credentials, or monoclaim credentials, represents the simplest selective disclosure technique. A multiclaim credential is partitioned into a subset of monoclaim credentials. Instead of having a single VC with $N$ claims, the holder will have $N$ VCs, each containing a single claim. Although this mechanism is compatible with existing solutions that do not yet support selective disclosure, it does not scale with credentials holding many data fields. Managing $N$ distinct credentials increases storage requirements, as each credential must include its metadata, cryptographic signatures, and auxiliary data. This also affects verification, as verifiers must individually validate each credential, raising computational and time costs. Furthermore, this approach incurs remarkable network overhead, as transmitting multiple credentials to disclose claims negatively impacts communication efficiency in bandwidth-constrained environments. These challenges are addressed by more efficient mechanisms, such as CSD-JWT, that do not need to generate multiple credentials to selectively disclose claims.
@@ VII-B Hashed Values
In hash-based methods, a VC includes the hashed values of each claim instead of the claim themselves. SD-JWT [13] is the state-of-the-art solution of JWT for selective disclosure. Each claim in the JWT payload is combined with a unique salt, hashed, and included in the credentials, ensuring claims are not directly exposed. The holder is given a SVC containing the plaintext claims and their corresponding salts. To disclose claims, the identity owner provides the verifier with the SD-JWT, the plain text claims they wish to disclose, and the corresponding salts. The verifier hashes each claim with its salt and checks if they match the pre-images of the hashed values in the SD-JWT. The main downside of this method is the credential size, which grows with the number of claims, impacting storage and transmission efficiency. Moreover, including hashes for all claims in the SD-JWT can make the system susceptible to inference attacks based on the credential’s structure. Similar to SD-JWT, De Salve et al. [52] propose using Hash-based Message Authentication Codes (HMACs) [53] to replace plaintext claims in VCs. The holder receives plain text claims and keys used to generate the HMACs. During the presentation, the holder discloses selected claims and keys, allowing the verifier to regenerate HMAC and validate HMAC. However, this approach shares the same limitations of SD-JWT. These drawbacks are addressed by CSD-JWT, which maintains a fixed credential size regardless of the number of included claims and prevents leakage of additional information (e.g., claim number).
Merkle trees [54] can also be used to create valid proofs for disclosed claims [55]. Each leaf node represents a claim’s hash, and parent nodes are obtained by hashing their children. Thus, knowing the root hash, a verifier can confirm a claim’s inclusion using a proof consisting of sibling hashes along the path to the root. This method ensures fast verification while keeping claims hidden but, contrary to CSD-JWT, results in variable proof lengths as the tree size grows with the number of claims in the credential. For binary trees, the path length is $log_{2}(N)$, where $N$ is the number of claims, potentially revealing a range for the VC’s field count.
@@ VII-C Signature Schemes
Signature schemes like Camenisch-Lysyanskaya (CL) signature [56] are widely used in Anonymous Credentials (ACs) [57], which can be seen as a special type of VCs, to ensure privacy guarantees such as anonymity or unlinkability. These schemes enable the holder to reveal specific claims while preserving anonymity. However, the applicability of ACs is limited being the authenticity of an identity or claim a primary requirement in most real-world scenarios [58].
Historically, the foundational work was done by the CL signature [56], which relies on the hard RSA assumption. Each message $m_{i}$ is mapped onto a finite field using modular exponentiation on random values $a_{i}$$b,s,c$ to generate the signature. Given the public key $pk=(n,a_{1},\dots,a_{n},b,c)$ and the signature $(v,e,s)$ the verification is performed as follows:
@@
For selective disclosure, the public key reveals $a_{i}^{m_{i}}$ rather than $a_{i}$. Several other schemes have been proposed [59, 60, 61, 62] with most of them using bilinear maps, allowing holders to derive valid signatures for subsets of signed messages. This prevents issuer-verifier collusion but introduces high computational and storage overhead especially on the holder side, making them less suitable for constrained environments. Such concerns do not affect CSD-JWT, as the holder is only required to select and sign the claim to disclose, without performing complex cryptographic operations.

### v2

@@ V-C Compromised Communication Channel
To prevent passive eavesdropping, secure communication among parties must be established. This can be effectively achieved by employing asymmetric encryption, utilizing the public keys stored in the DID Document of participants. Let $\mathcal{A}$ be an adversary attempting to tamper with the accumulator value $a$ or $WVC$. Any modification would require generating a valid witness for the altered data, which is infeasible without the issuer’s private key $sk_{i}$ under the EUF-CMA assumption. This ensures that any tampering with credential components is detectable and rejected by the verifier. CSD-JWT guarantees both binding and hiding properties, preserving both integrity and privacy.
@@ VI Performance Evaluation
Experiments were conducted by comparing CSD-JWT to SD-JWT and run on a 14th Gen Intel(R) Core(TM) i9-14900k that features 8 p-cores and 16 e-cores, and 48GB of RAM. For holder-side operations, we utilized the Nordic nRF52840 [nordic_nrf52840], an RFC7228 class-2 constrained device [rfc7228]. The nRF52840 is equipped with a 32-bit ARM Cortex-M4 processor, 1 MB of Flash memory, and 256 KB of RAM. We compared CSD-JWT to SD-JWT as it represents the current state-of-the-art solution and remains the only fully implemented and widely recognized solution available in the field at the time of this work. The experimental findings demonstrate that our mechanism is a valuable solution to enable selective disclosure while enhancing holder privacy compared to SD-JWT. For fairness, all comparisons between CSD-JWT and SD-JWT were conducted using the same claims within the credentials. All experiments were repeated $100$ times, and the results were averaged. A more extensive performance evaluation focusing also on Merkle Tree (MT) and BBS+ mechanisms for selective disclosure is provided in Appendix B.
(a) Issuer overhead (ms) for generating VCs.
(b) Storage requirements (KB) for VCs.
Figure 2: Performance metrics for VC issuance and storage.
@@ VI-A Implementation Setup
We implemented a prototype of CSD-JWT [^1] https://github.com/csdjwt/csd_jwt [^2] https://github.com/csdjwt/csd_jwt_artifact_evaluation and evaluated its performance while delivering the core functionalities. To effectively build the SD-JWT mechanism, we relied mainly on two libraries: josekit [^3] https://crates.io/crates/josekit, an OpenSSL-based library that provides ease of access to cryptographic signatures, and serde_json [^4] https://crates.io/crates/serde_json, a library that provides APIs to serialize and deserialize Rust objects to a JSON format. For generating accumulator values and the witnesses associated with claims, we relied on vb-accumulator [^5] https://crates.io/crates/vb-accumulator, developed by DockNetwork, which meets the requirements of [vitto_biryukov]. The underlying cryptographic accumulator is the ECC-based construction proposed in [vitto_biryukov], which supports efficient batch operations for both element accumulation and witness generation. The accumulator value is represented as a 256-bit elliptic curve point over the BN254 curve. To the best of our knowledge, this choice offers an optimal tradeoff among security guarantees, compact accumulator and witness sizes, and computational efficiency, making it particularly suitable for constrained environments. To evaluate the feasibility of CSD-JWT and SD-JWT on constrained holders, we developed a framework [^6] https://github.com/csdjwt/csd_jwt_constrained in C using the nRF Connect SDK and Parson [^7] https://github.com/kgabis/parson, a lightweight library for managing JSON data structures. Our credentials can include a variable number of claims, each formatted as claim_key:claim_value. For our evaluation, we consider VCs with up to 100 claims, reflecting variations across application domains [EU_Digital_Identity_Wallet_Use_Cases]. For instance, a VC representing an identification document typically contains around 15 claims. In contrast, in healthcare scenarios, the number of claims can vary significantly, ranging from approximately 10 claims, as in the case of the healthcare passport established by many countries for COVID-19 [covid19], to over 100 in detailed medical reports [THOMASON2021165]. Finally, for transparency and reproducibility, we open-sourced the proposed implementations.
VI-B Credential Issuance
Figure 2(a) compares the issuance latencies, expressed on a logarithmic scale, of the considered methods. SD-JWT achieves slightly better performance, mainly due to the CPU instructions used by the two mechanisms. Indeed, despite the great improvements in the instruction set of processors regarding finite field operations, hashing algorithms remarkably outperform them. However, it is worth noting that CSD-JWT only requires a few milliseconds even for credentials including 100 claims. Moreover, this does not represent a serious concern since, in real-world scenarios, issuers typically have powerful computing resources to manage the issuance process, significantly reducing these latencies.
VI-C Credential Storage
As shown in Figure 2(b), CSD-JWT remarkably outperforms SD-JWT even with a low number of claims in the VC. Using identical claims for both mechanisms, the primary difference lies in the size of the credentials and the storage of witnesses versus salts. The storage requirement for our method is much lower because CSD-JWT only contains the accumulator value, which is 256 bits. In contrast, SD-JWT comprises the hashes of the $N$ claims, i.e., $N\times$ 256 bits, making it dependent on the number of attributes. On the other hand, salts used in SD-JWT only require $N\times$ 128 bits, against the $N\times$ 256 bits for witnesses. Therefore, SD-JWT brings an overhead of ($N-2)\times 256$ bits for storing a VC with the same claims. As highlighted in the figure, CSD-JWT saves a significantly larger portion of memory while increasing the number of claims, achieving up to $46\%$ storage reduction. These results demonstrate that CSD-JWT is a valuable solution for selectively disclosing credentials in constrained environments. Moreover, this saving is even more impactful as each holder is expected to maintain multiple credentials [gartner2024digitalidentity, dayaratne2024ssi4iot].
@@ VI-D Generating and Verifying Verifiable Presentations
Figure 4: VP size reduction of CSD-JWT.
Size of Verifiable Presentations. Figures 3(b) and 3(c) report the results for VCs containing 10 and 100 claims, respectively. As clearly shown, CSD-JWT consistently generates smaller VPs than SD-JWT regardless of the number of disclosed claims. In Figure 3(b), when the holder discloses a single claim included in a VC with 10 claims, CSD-JWT achieves a $67\%$ reduction in the VP size, which decreases to $30\%$ when all the claims included in the VC are disclosed. The savings in the VP size are even more significant when considering VCs including a larger number of claims. Figure 3(c) demonstrates that with CSD-JWT, a VP disclosing 1 claim out of 100 is $93\%$ smaller in size than the same VP produced with SD-JWT. Even when all claims are being disclosed, CSD-JWT still provides a $27\%$ size reduction. Figure 4 illustrates the reduction in VP size by varying the rounded percentage of disclosed claims. These results highlight that CSD-JWT is an effective method to generate VPs with reduced size and network overhead, a feature that is particularly important in scenarios where holders’ devices have limited memory and network capabilities.
Figure 5: VP verification latency for a VC storing 100 claims.
[…]
@@ VI-E Discussion of Results
Credential Issuance. The time required to issue credentials based on CSD-JWT is higher than that for SD-JWT credentials, due to the increased complexity of the operations involved. In particular, the most demanding operation is the witness generation that, as demonstrated in [camacho2010impossibility], for $w$ witnesses demands at least $\Omega(w)$. However, the issuance time is a few milliseconds and does not represent a concern, as trusted authorities acting as issuers may have more powerful computing resources than those used for our experiments. Furthermore, it is worth noting that as these operations are always done on the issuer side, they do not affect at all the computing capabilities of the holder.
Credential Storage. In CSD-JWT, the storage requirement for the accumulator value remains constant, regardless of the number of claims included. This contrasts with SD-JWT, where the signed list contain a salted hash for each claim. Although salts are smaller in size compared to the witnesses, CSD-JWT offers an overall size reduction ratio of $27-93\%$ for VCs with 100 claims, with even greater savings for credentials containing more fields. In constrained environments, resources are shared across different functions. For instance, it is essential that the memory securely stores VCs and firmware, making even KB-level reductions valuable. Additionally, minimizing resource usage improves energy efficiency, a critical consideration in resource-limited settings. Finally, while our evaluation focuses on the overhead reduction for a single VC, identity owners are projected to use multiple VCs by 2026 [gartner2024digitalidentity], making the reductions offered by CSD-JWT significantly more remarkable.
Verifiable Presentation. The generation of VPs must be performed directly by holders. This is a logical consequence of the fact that they have to prove ownership over the disclosed information through a secret key. However, creating a VP does not introduce a particular overhead. Similar to SD-JWT, the holder is only required to create and sign a JWT containing the claims to be disclosed. This makes both CSD-JWT and SD-JWT practical on constrained devices.
On the other hand, we observed remarkable differences in size of the respective VPs. In SD-JWT, the credential shared during the presentation comprises the hashes of all the claims; thus, the size depends on the number of data included. In contrast, CSD-JWT maintains a constant-size accumulator value regardless of the number of claims, ensuring compactness. The only component that grows linearly with the number of disclosed claims is the set of witnesses required for verification. This minimizes the information shared, significantly reducing the network overhead and offering higher privacy.
@@ VII Related Work
This section reviews the primary selective disclosure mechanisms for VCs, categorizing them based on the underlying cryptographic techniques used to disclose claims. The categories include atomic credentials, hashed values, signature schemes, and Zero-Knowledge Proofs (ZKPs) [ETSI].
@@ VII-A Atomic Credentials
Atomic credentials, or monoclaim credentials, represents the simplest selective disclosure technique. A multiclaim credential is partitioned into a subset of monoclaim credentials. Instead of having a single VC with $N$ claims, the holder will have $N$ VCs, each containing a single claim. Although this mechanism is compatible with existing solutions that do not yet support selective disclosure, it does not scale with credentials holding many data fields. Managing $N$ distinct credentials increases storage requirements, as each credential must include its metadata, cryptographic signatures, and auxiliary data. This also affects verification, as verifiers must individually validate each credential, raising computational and time costs. Furthermore, this approach incurs remarkable network overhead, as transmitting multiple credentials to disclose claims negatively impacts communication efficiency in bandwidth-constrained environments. These challenges are addressed by more efficient mechanisms that do not need to generate multiple credentials to selectively disclose claims.
@@ VII-B Hashed Values
In hash-based methods, the claims in a VC are replaced with their salted hashes. SD-JWT [SD-JWT] is the state-of-the-art solution of JWT for selective disclosure, where each claim in the JWT payload is combined with a unique salt, hashed, and included in the credentials, ensuring claims are not directly exposed. The holder is given an SVC, containing the plaintext claims and their corresponding salts. To disclose claims, the identity owner provides the verifier with the SD-JWT, the plain text claims they wish to disclose, and the corresponding salts. The verifier hashes each claim with its salt and checks if they match the pre-images of the hashed values in the SD-JWT. The main downside of this method is the credential size, which grows with the number of claims, impacting storage and transmission efficiency. Moreover, including hashes for all claims in the SD-JWT can make the system susceptible to inference attacks based on the credential’s structure. Similar to SD-JWT, De Salve et al. [desalve] propose using Hash-based Message Authentication Codes (HMACs) [turner2008keyed] to replace plaintext claims in VCs. The holder receives plain text claims and keys used to generate the HMACs. During the presentation, the holder discloses selected claims and keys, allowing the verifier to regenerate HMAC and validate HMAC. However, this approach shares the same limitations as SD-JWT. These drawbacks are addressed by CSD-JWT, which prevents leakage of additional information (e.g., claim number) and substitutes the list of salted hashes with an accumulator value, whose size is constant regardless of the number of included claims.
Merkle Trees [merkle_tree_definition] can also be used to create valid proofs for disclosed claims [sd_merkle_tree]. Each leaf node represents a salted claim’s hash, and parent nodes are obtained by hashing their children. Thus, knowing the root hash, a verifier can confirm a claim’s inclusion using a proof consisting of sibling hashes along the path to the root. This method ensures fast verification while keeping claims hidden. However, unlike CSD-JWT, it results in variable proof lengths as the tree size grows with the number of claims in the credential. For binary trees, the path length is $log_{2}(N)$, where $N$ is the number of claims, potentially revealing a range for the VC’s field count.
@@ VII-C Signature Schemes
Signature schemes like Camenisch-Lysyanskaya (CL) signature [cl_sig_1] are widely used in Anonymous Credentials (ACs) [kakvi2023sok], which can be seen as a special type of VCs, to ensure privacy guarantees such as anonymity or unlinkability. These schemes enable the holder to reveal specific claims while preserving anonymity. However, the applicability of ACs is limited being the authenticity of an identity or claim a primary requirement in most real-world scenarios [vc_use_case].
Historically, the foundational work was done by the CL signature [cl_sig_1], which relies on the hard RSA assumption. Each message $m_{i}$ is mapped onto a finite field using modular exponentiation on random values $a_{i},b,s,c$ to generate the signature. Given the public key $pk=(n,a_{1},\dots,a_{n},b,c)$ and the signature $(v,e,s)$ the verification is performed as follows:
@@
For selective disclosure, the public key reveals $a_{i}^{m_{i}}$ rather than $a_{i}$. Several other schemes have been proposed [bbs, bbs_plus, cl_sig_2, idemix] with most of them using bilinear maps, allowing holders to derive valid signatures for subsets of signed messages. This prevents issuer-verifier collusion but introduces high computational and storage overhead, especially on the holder side, making them less suitable for constrained environments. Such concerns do not affect CSD-JWT, as the holder is only required to select and sign the VP, without performing complex cryptographic operations.